Assume that the significance level is a=0.1. Use the given information to find the P-value and the critical values.
The test statistic of z= 1.26 is obtained when testing the claim that p>0.9.
Answer #1

Given that, significance level (alpha) = 0.1

The null and the alternative hypotheses are,

0 : p0.9

Ha p0.9

Test statistic of z = 1.26

since, it is right tailed test,

p-value = P(Z > 1.26) = 1 - P(Z < 1.26) = 1 - 0.8962 = 0.1038

p-value = 0.1038

critical value at a = 0.1 is, z* = 1.28